# First Comparative Exoplanetology Within a Transiting Multi-planet System: Comparing the atmospheres of V1298 Tau b and c

### v1

@@ Appendix C Modelling the baseline
Young stars exhibit quasi-periodic flux modulation due to rotational variability (e.g see Feinstein et al. 2022; David et al. 2019b; Mann et al. 2016, 2022). The transit of planet c analyzed in this work occurred near an inflection point of the stellar baseline, resulting in significant deviation from a linear function. We modeled the baseline using linear, quadratic, and cubic polynomial functions for all the spectroscopic channels. The transit depths from all three fits were consistent within 1, but quadratic and cubic polynomials were statistically favored over a linear baseline (See Supplementary Figure 7).
@@
Appendix D Atmospheric models with ATMO
ATMO’s python based retrieval framework was applied and bench-marked in Lewis et al. (2020) and Rathcke et al. (2021) for HAT-P-41b and WASP-79b, respectively . This same retrieval framework is applied to V1298 Tau b and c in this work. This framework includes a Nested sampling Bayesian sampler DYNESTY (Speagle 2020) coupled to ATMO. We perform chemical equilibrium retrieval with isothermal - profile. Chemical equilibrium abundances are computed using Gibbs energy minimization as detailed in Drummond et al. (2016). We include H2-H2, H2-He Collision Induced opacity (CIA), as well as H2O, CH4 and CO high temperature molecular opacity in the retrieval. The clouds are treated as grey cloud deck between 1 and 50 millibar pressure levels, but with variable cloud opacity governed by the cloud factor used as the variable in the retrieval, as detailed in Goyal et al. (2018).
Appendix E Comparing V1298 Tau b ATMO retrieval with Barat et al. (2024)
For this study, we performed atmospheric retrieval on V1298 Tau b using the same setup as for planet c, assuming an isothermal atmosphere. Our results, indicating sub-solar metallicity and a mass upper limit of 20M⊕, align with Barat et al. (2024), who used PetitRadtrans. However, our retrieved temperatures are higher (800-900K) compared to their 600-700K. The discrepancy arises because Barat et al. (2024) used a non-isothermal temperature-pressure (T-P) profile with high interior temperature and vertical mixing to explain the non-detection of methane for this planet. Our isothermal model lacks this flexibility, leading to higher temperatures beyond the CH4/CO transition (Fortney et al. 2020) to eliminate methane from the upper atmosphere.
Appendix F Haze formation models
The haze microphysics models were simulated using the models described in Kawashima & Ikoma (2018), following the approach of Kawashima & Ikoma (2019). First, using a photochemical model the abundance of gaseous species was calculated. The XUV spectrum of V1298 Tau was adopted from Duvvuri et al. (2023). Haze production rates (or efficiency of haze formation) are uncertain, therefore we assumed a certain fraction of precursor molecules to be dissociated to form haze monomers. The haze precursor molecules are assumed to be CH4, HCN and C2H2 in our model. We derived vertical density and radius distributions of the haze particles following Kawashima & Ikoma (2018). We assumed solar elemental abundance ratio and other parameters such as eddy diffusion coefficient, monomer radius, internal density, and refractive index of haze particles similar to Kreidberg et al. (2022). We simulated transmission spectra, for V1298 Tau b and c assuming 23 and 20M⊕ respectively with obtained profiles of gas and haze particles. In Figure 1 we show the simulated transmission spectra assuming a 1 haze production efficiency.
Appendix G Theoretical evolutionary models
@@
Our models start from the time of disk dispersal and calculate the thermal evolution and structural evolution. We consider both core-envelope as well as gradually mixed interior structures. We include atmospheric mas loss due to photoevaporation from the topmost layer of the planets using the formalism of Rogers & Owen (2021). The mass loss rates are estimated from an energy-limited model (Erkaev et al. 2007), assuming a fixed photoevaporation efficiency. Our models include core erosion and mixing with the envelope. The evolutionary models are described in Vazan et al. (2022). We choose the initial conditions (mass, radius) such that it matches the current mass and radius for the planets. We assumed 23M⊕ and 20M⊕ for the current mass estimates of the V1298 Tau b and c respectively. We run simulations with both core-envelope and gradually mixed interior models. Both interior structure models can explain the current mass/radius of the planets. Mass and radius (normalized at current value) evolution models for the V1298 Tau planets are shown in Figure 2.
Appendix H Effect of stellar activity on transmission spectrum
@@
We estimate the effect of stellar activity on the transmission spectrum of V1298 Tau c. We adopt the method outlined in Rackham et al. (2019), where we estimate a correction factor based on the spot coverage fraction and spot temperature contrast. We adopt a spot coverage fraction of 20% for V1298 Tau (Feinstein et al. 2021). We adopt an extreme spot contrast of 1000K following spot contrasts derived for a sample of T Tauri stars (Koen 2016). We show a comparison between the observed and corrected transmission spectrum in Supplementary Figure 10. The observed and contamination corrected normalized spectra are consistent within 1 for all spectroscopic channels.
@@
Since V1298 Tau is a K-type star, its contamination function does not show molecular features like M-dwarfs (Barclay et al. 2021). Therefore, the effect of the stellar correction is to change the level of the continuum in this case. However, the atmospheric features, which are relative to the continuum remain largely unaffected.
Appendix I Supplementary Figures
@@
| Planet name | Mass [M⊕] | Atmospheric metallicity [solar] | Semi-major axis [AU] | Stellar metallicity [dex] | Reference |
| Wasp-18b | 10.200.35 | 1.03 | 0.02 | 0.110.08 | Coulombe et al. (2023) |
| Wasp-39b | 0.2800.03 | 10 | 0.048 | -0.030.1 | Rustamkulov et al. (2023) |
| Wasp-77Ab | 1.660.06 | 0.12 | 0.023 | -0.010.1 | August et al. (2023) |
| Wasp-96b | 0.480.03 | 7.5 | 0.045 | 0.140.19 | Radica et al. (2023) |
| TOI-270d | 0.0150.001 | 225 | 0.07 | -0.20.1 | Benneke et al. (2024) |
| Wasp-107b | 0.10.005 | 10 | 0.05 | 0.020.09 | Welbanks et al. (2024) |
| HD149026b | 0.3580.01 | 20 | 0.0436 | 0.330.1 | Gagnebin et al. (2024) |
| HD209458b | 0.59029040.01 | 3 | 0.05 | -0.010.05 | Xue et al. (2024) |
| Wasp-80b | 0.530.03 | 6 | 0.03 | -0.10.1 | Bell et al. (2023) |
| Wasp-17b | 0.780.2 | 50 | 0.051 | -0.070.1 | Grant et al. (2023) |
| GJ1214b | 0.0260.001 | 100 | 0.01 | 0.020.1 | Kempton et al. (2023) |

### v2

@@ Appendix C Modeling the baseline
Young stars exhibit quasi-periodic flux modulation due to rotational variability (e.g. see David et al. 2019b; Mann et al. 2016, 2022). The transit of planet c analyzed in this work occurred near an inflection point of the stellar baseline, resulting in significant deviation from a linear function. We modelled the baseline using linear, quadratic, and cubic polynomial functions for all the spectroscopic channels. The transit depths from all three fits were consistent within 1, but quadratic and cubic polynomials were statistically favored over a linear baseline (See Fig. 8).
@@
Appendix D Theoretical evolutionary models
@@
The evolutionary models we use in the current study are described in Vazan et al. (2022). Our models start from the time of disk dispersal and calculate the thermal evolution and structural evolution. We consider both core-envelope as well as gradually mixed interior structures. We include atmospheric mas loss due to photoevaporation from the topmost layer of the planets using the formalism of Rogers & Owen (2021). The mass loss rates are estimated from an energy-limited model (Erkaev et al. 2007), assuming a fixed photoevaporation efficiency. Our models include core erosion and mixing with the envelope. We choose the initial conditions (mass, radius) such that it matches the current mass and radius for the planets. We assumed 17 M⊕ (peak of ATMO retrieval posterior distribution) for the current mass estimates of the V1298 Tau c. For planet b, we show two models: 10 M⊕ (consistent within 1 to median mass retrieved from ATMO) and 23 M⊕ (3 mass upper limit from Barat et al. (2024)). We run simulations with both core-envelope and gradually mixed interior models. Both interior structure models can explain the current mass/radius of the planets. Mass and radius (normalized at current value) evolution models for the V1298 Tau planets are shown in Figure 4.
Appendix E Effect of stellar activity on transmission spectrum
@@
We estimate the effect of stellar activity on the transmission spectrum of V1298 Tau c. We adopt the method outlined in Rackham et al. (2019), where we estimate a correction factor based on the spot coverage fraction and spot temperature contrast. We adopt a spot coverage fraction of 20% for V1298 Tau (Feinstein et al. 2021). We adopt an extreme spot contrast of 1000 K following spot contrasts derived for a sample of T Tauri stars (Koen 2016). We show a comparison between the observed and corrected transmission spectrum in Fig. 11. The observed and contamination corrected normalized spectra are consistent within 1 for all spectroscopic channels.
@@
Since V1298 Tau is a K-type star, its contamination function does not show molecular features like M-dwarfs (Barclay et al. 2021). Therefore, the main effect of the stellar correction is to change the level of the continuum in this case. However, the atmospheric features, which are relative to the continuum remain largely unaffected.
Appendix F Supplementary Figures
@@
| Planet name | Mass [MJ] | Atmospheric metallicity [solar] | Semi-major axis [AU] | Stellar metallicity [dex] |
| WASP-18b 11111111Coulombe et al. (2023) | 10.200.35 | 1.03 | 0.02 | 0.110.08 |
| WASP-39b 2,3,4,523452,3,4,52 , 3 , 4 , 52,3,4,523452,3,4,52 , 3 , 4 , 5Rustamkulov et al. (2023); Feinstein et al. (2023); Ahrer et al. (2023); Alderson et al. (2023) | 0.2800.03 | 10 | 0.048 | -0.030.1 |
| WASP-77Ab 66666666August et al. (2023) | 1.660.06 | 0.12 | 0.023 | -0.010.1 |
| WASP-96b 77777777Radica et al. (2023) | 0.480.03 | 0.23 | 0.045 | 0.140.19 |
| TOI-270d 88888888Benneke et al. (2024) | 0.0150.001 | 225 | 0.07 | -0.20.1 |
| WASP-107b 99999999Sing et al. (2024) | 0.10.005 | 43 | 0.05 | 0.020.09 |
| HD149026b 1010101010101010Gagnebin et al. (2024) | 0.3580.01 | 20 | 0.0436 | 0.330.1 |
| HD209458b 1111111111111111Xue et al. (2024) | 0.59029040.01 | 3 | 0.05 | -0.010.05 |
| WASP-80b 1212121212121212Bell et al. (2023) | 0.530.03 | 6 | 0.03 | -0.10.1 |
| WASP-17b 1313131313131313Grant et al. (2023) | 0.780.2 | 50 | 0.051 | -0.070.1 |
| GJ1214b 1414141414141414Kempton et al. (2023) | 0.0260.001 | 100 | 0.01 | 0.020.1 |